OpenStudy (anonymous):

True of false? Sin 2x = 2sin x cos x

OpenStudy (anonymous):

Are they asking for proof? Use the sum and difference formula for sin(u+v)=sinucosv+cosusinv 2x=x+x, now sub x for u and x for v

OpenStudy (anonymous):

sin(x+x)=sinxcosx+cosxsinx sin(2x)= 2cosxsinx

OpenStudy (anonymous):

thanks this statement is true